OpenStudy (anonymous):

Find the slope of the tangent line to the graph of y=x+cos(xy) at x=0

OpenStudy (anonymous):

Can u find the derivative of the above expression?

OpenStudy (anonymous):

yes? 1-siny?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

It is\[\frac{ dy }{ dx }=1-\sin(xy)(y+x \frac{ dy }{ dx })\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

oh I see! the chain rule?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Ok what is the slope then?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

Nope it is 1

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Insert 0 for x and 1 for y

OpenStudy (anonymous):

ohh okay I see thank you!